Given the following functions; 

TC= 50Q+20Q2(Cost Function) 

P= 400-20Q (Demand function) 

Determine the profit maximizing level of price and output. 


1
Expert's answer
2021-08-27T11:57:52-0400

"TC=50Q+20Q^2"

"MC=50+40Q"

"P=400-20Q"

For any linear demand curve of the form "P=a-bQ" , the marginal revenue function has the form "MR=a-2bQ"

"\implies MR=400-2(20)Q"

"MR=400-40Q"

At maximum profit, "MC=MR"

"50+40Q=400-40Q"

"Q=4.375"

"P=400-20Q"

"P=400-20(4.375)"

"P=400-87.5"

"P=312.5"

The profit maximizing level of price is 312.5 and output is 4.375